$SPAD/src/input coercels.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool coercels.output
)set message test on
)set message auto off
)clear all
--S 1 of 8
alternatingGroup 4
--R
--R
--R   (1)  <(1 2)(3 4),(1 2 3)>
--R                                              Type: PermutationGroup(Integer)
--E 1
--S 2 of 8
% :: List Permutation Integer
--R
--R
--R   (2)  [(1 2)(3 4),(1 2 3)]
--R                                             Type: List(Permutation(Integer))
--E 2
--S 3 of 8
li := %
--R
--R
--R   (3)  [(1 2)(3 4),(1 2 3)]
--R                                             Type: List(Permutation(Integer))
--E 3
--S 4 of 8
pgr := MonoidRing(Polynomial PrimeField 5, Permutation Integer)
--R
--R
--R   (4)  MonoidRing(Polynomial(PrimeField(5)),Permutation(Integer))
--R                                                                 Type: Domain
--E 4
--S 5 of 8
p : pgr := first  li

--R
--R
--R   (5)  (1 2)(3 4)
--R             Type: MonoidRing(Polynomial(PrimeField(5)),Permutation(Integer))
--E 5
--S 6 of 8
q : pgr := first  li
--R
--R
--R   (6)  (1 2)(3 4)
--R             Type: MonoidRing(Polynomial(PrimeField(5)),Permutation(Integer))
--E 6
--S 7 of 8
basis  := [p,q,p*p,p*q, q*p,q*q, p*q*q, p*q*p, q*p*q,q*q*p,q*p*q*q,q*q*p*q]
--R
--R
--R   (7)
--R   [(1 2)(3 4), (1 2)(3 4), 1, 1, 1, 1, (1 2)(3 4), (1 2)(3 4), (1 2)(3 4),
--R    (1 2)(3 4), 1, 1]
--R       Type: List(MonoidRing(Polynomial(PrimeField(5)),Permutation(Integer)))
--E 7
--S 8 of 8
% :: Set          MonoidRing(Polynomial PrimeField 5,Permutation Integer)
--R
--R
--R   (8)  {(1 2)(3 4),1}
--R        Type: Set(MonoidRing(Polynomial(PrimeField(5)),Permutation(Integer)))
--E 8
)spool
)lisp (bye)

———-

References^{[1] nothing}

5